Consider the following table:

Stock Fund Bond Fund
Scenario Probability Rate of Return Rate of Return
Severe recession 0.10 −43% −12%
Mild recession 0.20 −17.0% 12%
Normal growth 0.30 17% 6%
Boom 0.40 31% 4%

a. Calculate the values of mean return and variance for the stock fund. (Do not round intermediate calculations. Round "Mean return" value to 1 decimal place and "Variance" to 4 decimal places.)

b.Calculate the value of the covariance between the stock and bond funds. (Negative value should be indicated by a minus sign. Do not round intermediate calculations. Round your answer to 4 decimal places.)
Answer #1
Scenario Probability Stock Fund return

Bond fund return

Severe recession 0.1 -43% -12%
Mild recession 0.2 -17% 12%
Normal growth 0.3 17% 6%
Boom 0.4 31% 4%

Stock Fund

Expected return is calculated using the formula:

Expected return = E[R] = p1*R1 + p2*R2 + p3*R3 + p4*R4

Expected return of stock Fund = E[RS] = 0.1*(-43%) + 0.2*(-17%) + 0.3*17% + 0.4*31% = -4.3% + (-3.4%) + 5.10% + 12.4% = 9.80%

Expected return or mean return of stock Fund = 9.8%

Variance is calculated using the formula:

Variance = p1*(R1 - E[R])2 + p2*(R2 - E[R])2 + p3*(R3 - E[R])2 + p4*(R4 - E[R])2

Variance of the stock Fund = σS2 = 0.1*(-43%-9.8%)2 + 0.2*(-17% - 9.8%)2 + 0.3*(17% - 9.8%)2 + 0.4*(31% - 9.8%)2 = 0.0278784 + 0.0143648 + 0.0015552 + 0.0179776 = 0.061776

We need to calculate the variance in %-squared. So to convert it into %-squared we need to multilply it by 10000.

Variance of the stock fund in %-squared = 0.061776*10000 = 617.76

Answer part a

Mean Return (%) = 9.8

Variance (%-Squared) = 617.76

Part b

Expected return of the bond fund = E[RB] = 0.1*(-12%) + 0.2*12% + 0.3*6% + 0.4*4% = -1.2% + 2.4% + 1.8% + 1.6% = 4.6%

We have the following data:

p1 = 0.1, p2 = 0.2, p3 = 0.3, p4 = 0.4

R1,S = -43%, R2,S = -17%, R3,S = 17%, R4,S = 31%

R1,B = -12%, R2,B = 12%, R3,B = 6%, R4,B = 4%

Expected return of the stock Fund = E[RS] = 9.8%, Expected return of the bond Fund = E[RB] = 4.6%

Covariance between the stock fund and the bond fund is calculated using the formula:

Covariance between stock fund and bond fund = Cov(S,B) = p1*(R1,S - E[RS])*(R1,B - E[RB]) + p2*(R2,S - E[RS])*(R2,B - E[RB]) + p3*(R3,S - E[RS])*(R3,B - E[RB]) + p4*(R4,S - E[RS])*(R4,B - E[RB])

Cov(S,B) = 0.1*(-43% - 9.8%)*(-12% - 4.6%) + 0.2*(-17% - 9.8%)*(12% - 4.6%) + 0.3*(17% - 9.8%)*(6% - 4.6%) + 0.4*(31% - 9.8%)*(4% - 4.6%) = 0.0087648 + (-0.0039664) + 0.0003024 + (-0.0005088) = 0.004592

To convert the covariance to %-squared, we need to multiply it by 10000

Cov(S,B) in %-squared = 0.004592*10000 = 45.92

Answer b -> Covariance (%-Squared) = 45.92